OpenStudy (anonymous):

If the APY of a savings account is 2.5%, and if the principal of the savings account was $2000 an entire year, waht will the balance of a savings account be if all the interest is paid for the year?

OpenStudy (anonymous):

i = prt i = 2000*.025*1 i = 50 Ending balance is p + i = 2000 + 50 = 2050